OpenStudy (anonymous):

What is the product of 5*sqrt(u)(sqrt(3) + sqrt(147) 5*sqrt(3u) + 7*sqrt(3) 40*sqrt(3u) 54*sqrt(3u) 35u*sqrt(3)

OpenStudy (anonymous):

I have no idea at all, please help me with the answer!

OpenStudy (rajee_sam):

\[5\sqrt{u}(\sqrt{3} +\sqrt{147})\]

OpenStudy (rajee_sam):

that is your problem right??

OpenStudy (anonymous):

Yes it is

OpenStudy (rajee_sam):

\[=5\sqrt{u} ( \sqrt{3} + \sqrt{49 \times 3})\]\[=5\sqrt{u} ( 1\sqrt{3} + 7\sqrt{3})\]\[= 5\sqrt{u} ( 8\sqrt{3})\]\[= 40 \sqrt{u} \times \sqrt{3}\]\[= 40 \sqrt{3u}\]

OpenStudy (anonymous):

Thank you so much!!